OpenStudy (aonz):

Easy question for most people :P WIll give medal and fan :) An object is brought to rest from a velocity of 24m/s due north in 16 seconds. Find (a) the displacement , (b) acceleration and distance travelled

OpenStudy (aonz):

for i think you use this formula am i correct? s= ut + 1/2at^2

OpenStudy (badhi):

since you know the starting velocity and end velocity as well as the time spent, you can use the following equations $$S=\frac{(u+v)}{2}t$$ and for acceleration, $$v=u+at$$

OpenStudy (aonz):

i just started physics and i dont know this formula yet :P S=(u+v) /2 * t

OpenStudy (aonz):

our teacher gave us 3 formulas to work with. 1-3 equations of motion

OpenStudy (anonymous):

yes threre are three equations \[\Large v_{f}=v_{i}+at\] \[\Large S=v_{i}t+1/2at^2\] \[\Large 2as=v_{f}^2-v_{i}^2\] where vf=final velocity vi=initil veloct s=distance covered a=acceleration t=time

OpenStudy (anonymous):

using the first equation since objecvt is brought to rest so vf=0 vi=24m/sec t=16 we can find a using first equation \[\Large 0=24+16a\] \[\Large a=-24/16=-1.5m/\sec^2\] use the other equation to get distance .

OpenStudy (aonz):

we use which equation? number 2?

OpenStudy (aonz):

im pretty sure thats not right... The object cannot be deccelerating???

OpenStudy (badhi):

If it is not deccelerating how can a moving object be brought to rest?

OpenStudy (aonz):

oh yes :P

OpenStudy (aonz):

so im going to use which formula to work out displacement? 2nd or 3rd I think 2nd

OpenStudy (anonymous):

the object is decelarating because you have brought it to rest . think of if you are driving a car and you apply brakes what will happen ???? acceleration or deceleration ??

OpenStudy (aonz):

deceleration yea i understand that now